You own $17,290 of Opsware, Inc. stock that has a beta of 3.61. You also own $20,020 of Lowe’s Companies (beta = 1.62) and $8,190 of New York Times (beta = 1.16). Assume that the market return will be 15 percent and the risk-free rate is 6 percent. What is the market risk premium? What is the risk premium of each stock? (Round your answers to 2 decimal places.) What is the risk premium of the portfolio? (Do not round intermediate calculations and round your final answer to 2 decimal places.) Next Visit question mapQuestion 9 of 11 Total 9 of 11 Prev
Answer #1

1.
=market return-risk free rate
=15%-6%=9%

2.
Risk premium=beta*Market Risk Premium
Opsware=3.61*9%=32.490%

Lowe's=1.62*9%=14.580%

New York Times=1.16*9%=10.440%